OpenStudy (roberts.spurs19):

The moment of inertia of a thin rod of length 2L and mass M about an axis normal to its length and acting through the rod at a distance 3L/4 from one of its ends is given by which one of the following and why? a)19ML^2/48 b)15ML^2/16 c)27ML^2/8 d)19ML^2/64 e)8ML^2/27 Thank you!

OpenStudy (irishboy123):

 this is based on the definition of inertia \[dI = r^2 dm\] the small element of the rod, length dr, as shown, has mass dm. the density per unit length of the rod is M/2L so for that small element we can say dm = M/2L dr, and as dI = r^2 dm , thus dI = M/2L r^2 dr which delivers this integral: \[I = \frac{M}{2L} \int\limits_{-\frac{3L}{4}}^{\frac{5L}{4}} r^2 dr = \frac{M}{6L} \left| r^3 \right|^{\frac{5L}{4}}_{-\frac{3L}{4}}\]